byeeee:

What is the value of x in the solution to the following system of equations? x − y = −3 x + 3y = 5 Answer for Blank 1:

byeeee:

@Vocaloid

Vocaloid:

since you have -y in the first equation and +3y in the second equation try multiplying the first equation by 3, adding the equations together, then solving for x

byeeee:

what do you mean by multiplying the first equation by three? @Vocaloid do you mean multiply 3 by -3?

Vocaloid:

you can multiply both sides of the equation by 3 3(x − y) = 3(−3) then distribute

byeeee:

okay

byeeee:

so i got -9

Vocaloid:

don't forget about the left side 3(x-y) what do you get when you distribute this?

byeeee:

you get 3x-3y

Vocaloid:

awesome so 3x - 3y = -9 combine that with our original equation to get the new system 3x - 3y = -9 x + 3y = 5 now try adding the two equations together

byeeee:

okay

byeeee:

im a tad confused

Vocaloid:

take it one set of variables at a time notice how one equation has x and the other has 3x so x + 3x = ?

byeeee:

5

Vocaloid:

hm, not quite. notice how we have one x + three x if 1 + 3 = 4 what might x + 3x = ?

byeeee:

im not sure this one is really confusing me im sorry hm

Vocaloid:

well think of the x's as just ordinary objects 1 object + 3 objects = 4 objects x + 3x = 4x

Vocaloid:

using that same logic, let's go onto the y terms any ideas what -3y + 3y = ?

byeeee:

0?

Vocaloid:

awesome so finally we have -9 + 5 which is just -4 so 4x = -4 any ideas what x might be equal to?

byeeee:

-1?

Vocaloid:

yup, so -1 is your solution

byeeee:

wow thank you so much! @Vocaloid